# Pitfalls of Rule- and Model-based Verifiers – A Case Study on Mathematical Reasoning

### v1

@@ Abstract
Trustworthy verifiers are essential for the success of reinforcement learning with verifiable reward (RLVR), which is the core methodology behind various large reasoning models such as DeepSeek-R1. In complex domains like mathematical reasoning, rule-based verifiers have been widely adopted in previous works to train strong reasoning models. However, the reliability of these verifiers and their impact on the RL training process remain poorly understood. In this work, we take mathematical reasoning as a case study and conduct a comprehensive analysis of various verifiers in both static evaluation and RL training scenarios. First, we find that current open-source rule-based verifiers often fail to recognize equivalent answers presented in different formats across multiple commonly used mathematical datasets, resulting in non-negligible false negative rates. This limitation adversely affects RL training performance and becomes more pronounced as the policy model gets stronger. Subsequently, we investigate model-based verifiers as a potential solution to address these limitations. While the static evaluation shows that model-based verifiers achieve significantly higher verification accuracy, further analysis and RL training results imply that they are highly susceptible to hacking, where they misclassify certain patterns in responses as correct (i.e., false positives). This vulnerability is exploited during policy model optimization, leading to artificially inflated rewards. Our findings underscore the unique risks inherent to both rule-based and model-based verifiers, aiming to offer valuable insights to develop more robust reward systems in reinforcement learning.
@@ 1 Introduction
Figure 1: The training and evaluation curves of RL on Qwen-2.5-7B using different verifiers, with the x-axis representing training iterations in all plots. Left illustrates the evaluation accuracy averaged over multiple benchmarks, including GSM8K, MATH500, Minerva Math, OlympiadBench, AIME24, and AMC23. Right depicts changes in reward values during training. The “training rewards” indicate the rewards provided by the corresponding reward system to the policy model, whereas the “oracle rewards” represent rewards the model receives when judged by combining with GPT-4o. We provide detailed breakdown of evaluation results in Table 2.
In our subsequent RL training experiments, however, we observe that model-based verifiers introduce unique challenges and yield mixed outcomes: while some verifiers can improve RL results by an average of more than 3 absolute points over rule-based verifiers, others are vulnerable to hacking, leading to suboptimal results of RL training(see the left side of Figure 1). Reward hacking – a well-known issue in RL – refers to the exploitation of specific patterns by the policy model to deceive the reward system and obtain artificially high rewards (illustrated in the bottom right of Figure 1). Notably, we find that our trained verifier, despite achieving higher classification accuracy than off-the-shelf alternatives, is more susceptible to hacking during RL training. These findings indicate that the classification accuracy of a verifier does not necessarily reflect its resistance to reward hacking, and therefore may not be a reliable indicator of its effectiveness in RL training.
In the final part of our study, we conduct a pilot investigation into specific patterns that can exploit vulnerabilities in verifiers. We construct a range of adversarial patterns inspired by our case studies, such as the insertion of empty characters or garbled text. Using these constructed “hacking data”, we evaluate whether various model-based verifiers can be deceived. Our results show that most verifiers are easily fooled by these patterns and the discriminative verifiers are more robust than the generative ones. While fine-tuning improves static evaluation scores, it does not necessarily enhance robustness and, in some cases, even worsens it – an observation consistent with our RL experimental findings.
Our findings in this work clearly identify the pitfalls of both rule-based and model-based verifiers in the context of mathematical reasoning: current rule-based verifiers are not sufficiently accurate even for widely used open-source mathematical datasets with short answers that should be easily verifiable. Pursuing more accurate model-based verifiers is a promising direction to improve RL performance; however, this approach introduces unique vulnerabilities to hacking, which require further investigation in future work.
@@ 3.1 Evaluation Dataset Construction
We curate our dataset as a static classification task to examine the capabilities of verifiers in classifying the correctness of model responses with respect to a provided ground-truth answer. The curation process involves three main steps: First, we select and sample from four mathematical RL datasets – Math (Hendrycks et al., 2021), DeepscaleR (Luo et al., 2025a), Open-Reasoner-Zero (ORZ-Math)(Hu et al., 2025), and Skywork-OR1(He et al., 2025) – with 1,000 queries sampled from each dataset. In the second step, we generate two responses for each of these queries using two types of language models: (1) Short-CoT models, specifically Qwen2.5-Math-7B-Instruct (Yang et al., 2024a) and Qwen2.5-32B-Instruct (Yang et al., 2024b), and (2) R1-style long CoT models, namely Deepseek-R1-Distill-Qwen-7B and 32B (DeepSeek-AI et al., 2025). Finally, we employ GPT-4o (Hurst et al., 2024) as an annotator to provide ground-truth annotations based on the response and target answer, on whether the model’s response aligns with the target answer, based on a prompt shown in Figure 4 in Appendix A. The final dataset comprises 2,000 examples per dataset, for a total of 8,000 examples. We emphasize that the datasets we selected already represent a relatively easy setting for verification – these datasets contain only short answers, and most were specifically curated to be easily verifiable by rules in order to facilitate RL. Consequently, more realistic scenarios are likely to present greater challenges than those reflected in our empirical results next.
Justification of GPT-4o annotation.
@@ 3.2 Rule-based Verifiers: Precision at the Cost of Recall
We adopt three popular rule-based verifier implementations including: (1) Verl Math Verifier, [^1] https://github.com/volcengine/verl/blob/main/verl/utils/reward_score/math.py (2) Qwen-Math Verifier, [^2] https://github.com/QwenLM/Qwen2.5-Math/blob/main/evaluation/evaluate.py and (3) HuggingFace Math Verifier, [^3] https://github.com/huggingface/Math-Verify which have been widely utilized in previous studies (Zeng et al., 2025b, a; He et al., 2025; Yu et al., 2025).
@@
To evaluate the performance of these verifiers, we test them on a subset of data sampled from Deepseek-R1-Distill-Qwen-32B, a state-of-the-art open-source model known for its exceptional mathematical reasoning abilities. As shown in Table 5 in Appendix B, all three verifiers exhibit near-perfect precision, consistently achieving over 99% precision. This means that if an answer passes the rules, it is almost certainly correct because the rule-based verifiers rely on deterministic programming language logic and computation. Notably, the HuggingFace Math Verifier and Qwen-Math Verifier show very similar performance. However, the rigid structure of these rule-based systems presents challenges when dealing with more complex or edge-case queries, leading to a significantly low recall rate of 0.78 on some datasets like Skywork-OR1, as shown in Figure 3. This indicates that there are some correct responses that are misjudged as incorrect, and we illustrate some cases in Figure 5.
@@
Figure 3 presents the recall rate of HuggingFace Math Verifier across various sampling models and datasets. A key observation is that as the capabilities of the models increase, providing accurate supervision becomes more challenging for rule-based verifiers. For example, the recall rate for the Long-CoT models, such as DeepSeek-R1-Distill-Qwen-7B and 32B averages around 0.92, which is much lower than other weaker models. This is because some complex queries, which only advanced models can solve, are misjudged by the rule-based verifier. The inability of rule-based verifiers underlines the difficulty in verifying highly capable models. This trend is particularly concerning, given that the community is advancing increasingly powerful reasoning models, which in turn require stronger verifiers.
@@
As shown in Figure 3, more challenging datasets tend to have lower recall rates. For example, the Math dataset, known for its simplicity and well-established structure, results in a relatively high recall rate, while newer and more difficult datasets, such as Skywork-OR1, experience significantly lower recall rates. These findings underscore a critical limitation: As datasets become more varied, and more challenging, the reliability of rule-based verifiers as supervision tools for scalable reinforcement learning diminishes.
Table 1: The performance of various model-based verifiers across different datasets is presented, with results shown in terms of Precision/Recall. To assess model-based verifiers within a hybrid verifier framework, we evaluate a subset sampled from DeepSeek-R1-Distill-Qwen-32B and exclude examples already verified as correct by the HuggingFace Math Verifier. As a result, the HuggingFace verifier shows zero metrics here, since it classifies all remaining examples as incorrect. “DS” denotes DeepSeek, and for Qwen series models, the “instruct” suffix is omitted for clarity.
@@ 3.3 Model-based Verifiers: Toward Greater Flexibility
To mitigate the limitation of rule-based verifiers, we next investigate model-based verifiers as a potential alternative. Model-based verifiers seek to leverage the core capabilities of LLMs, including their advanced reasoning skills, to produce more accurate judgments. They are, in principle, better equipped to evaluate answers presented in diverse formats. Model-based verifiers are explored in several concurrent works (Su et al., 2025; Ma et al., 2025; Seed et al., 2025) without deep discussion or ablation on their strengths and limitations. As such, the potential benefits of model-based verifiers over rule-based ones in the context of RL remain unclear. Additionally, model-based verifiers, especially generative ones, introduce significant computational overhead, which can be critical for online RL training. In this section, we first explore model-based verifiers in static evaluation, and in §4 we will discuss its effect in RL training.
@@
We evaluate two categories of general LLM as a verifier: (1) Short-CoT models: Qwen2.5-instruct (1.5B and 7B) (Yang et al., 2024b) and Qwen2.5-Math-instruct (1.5B and 7B) (Yang et al., 2024a). (2) R1-style long-CoT models: DeepSeek-R1-Distill-Qwen (1.5B and 7B) (DeepSeek-AI et al., 2025). We will also discuss model-based verifiers specifically trained for verification tasks in §5. Note that we focus on models with up to 7B parameters, as larger models are neither practical nor efficient for scaling RL training. We note that all these models are generative which will typically generate reasoning traces along with the final judgment. Since rule-based verifiers achieve nearly perfect precision but tend to produce false negatives, we focus here exclusively on the examples that rule-based verifiers classify as incorrect. This approach is able to better distinguish different model-based verifiers. It also aligns with the design of our hybrid verification system in the RL experiments, where rule-based verifiers are applied first, and model-based verifiers are used only for those cases deemed incorrect. We will provide further details in §4.1. Specifically, for the evaluation dataset, we use the subset sampled from DeepSeek-R1-Distill-Qwen-32B, excluding examples that have already been classified as correct by the HuggingFace Math Verifier. For additional details about the evaluation procedure, please refer to Appendix C.
@@ 4 The Effect of Verifiers on RL Training
In §3, we showed that model-based verifiers achieve strong performance across datasets and substantially improve recall on the verification task. Building on this, we adopt model-based verifiers in RL training and compare their impact with rule-based verifiers. Specifically, we propose a hybrid verifier that integrates the strengths of both approaches. We first evaluate its performance in static settings, then analyze its improvements over rule-based verifiers in RL training, as well as its training efficiency compared to fully model-based verifiers.
@@ 4.1 The Hybrid Verifier
In Table 6 in Appendix D, we present the static evaluation results of the rule-based verifiers, the model-based verifiers, and the hybrid verifiers. The hybrid verifier improves recall by approximately 3 percentage points on average over the rule-based verifier, while consistently maintaining precision above 98%. Model-based verifiers alone may exhibit lower recall than the hybrid approach, as smaller models can overthink some straightforward cases. However, integrating the rule-based verifier mitigates this issue, resulting in superior overall performance. In general, the hybrid system achieves superior performance in both precision and recall. Furthermore, by filtering out straightforward cases that the rule-based verifier can confidently resolve, the hybrid design substantially reduces the computational load on the model-based verifier. We discuss this further in §4.3.
@@ 4.2 Experimental Setup
For all experiments, we follow the approach of Deepseek-R1 (DeepSeek-AI et al., 2025), using GRPO (Shao et al., 2024) as the training algorithm and adhering to the zero RL training recipe — starting training directly from the base model. Our policy model is Qwen2.5-7B Base (Yang et al., 2024b). We conduct RL training using the DeepscaleR dataset and construct a hybrid verifier by combining the Huggingface Math Verifier with DeepSeek-R1-Distill-Qwen-1.5B, the best-performing 1.5B model on DeepscaleR, as shown in Table 1. Further training details are provided in Appendix E
@@
We build our evaluation script based on  Yang et al. (2024a), which utilize the rule-based verifier. And we evaluate performance on standard mathematical reasoning benchmarks, including GSM8K (Cobbe et al., 2021), MATH 500 (Hendrycks et al., 2021), OlympiadBench (He et al., 2024) and Minerva Math (Lewkowycz et al., 2022), as well as on competition-level benchmarks including AIME 2024 and AMC 2023. Following the approach in (Zeng et al., 2025a; Yu et al., 2025), we evaluate AIME 2024 by averaging the results of 32 random samplings (Avg@32) to ensure stable evaluation. For additional evaluation details, please refer to Appendix E.
@@ 4.3 Results
Table 2: Detailed performance of models across multiple benchmarks. The best result from each run is reported. Blue lines indicate models trained with a hybrid verifier without evidence of reward hacking, while pink lines indicate runs where reward hacking is detected. “HF” represents HuggingFace Math Verifier. Training and evaluation curves for these models are presented in Figure 1 and Figure 8.
@@
AMC23 Avg.
Qwen2.5-7B-SimpleRL-Zoo 91.7 78.2 38.6 40.4 15.6 62.5 54.5
Qwen2.5-7B 88.2 64.6 25.7 30.1 0.3 30.0 39.8
      $\hookrightarrow$ + DeepscaleR & HF verifier 92.8 80.0 37.5 42.2 15.3 62.5 55.1
          $\hookrightarrow$ + DS-R1-Distill-Qwen-1.5B verifier 93.3 82.4 41.2 42.5 20.4 70.0 58.3
          $\hookrightarrow$ + R1-Distill-Verifier-1.5B verifier 93.0 79.8 40.4 40.1 17.8 77.5 58.1
          $\hookrightarrow$ + general-verifier 92.5 82.0 43.0 40.9 18.4 70.0 57.8
@@
As shown in Figure 1, the inclusion of the hybrid verifier significantly improves the accuracy of evaluation with the peek point of 58.35, more than 3 points higher than that of using rule-based verifier only. Moreover, the hybrid verifier boosts dataset utilization by reducing the proportion of answers that cannot be correctly parsed. For instance, as shown in Table 2, the performance of the rule-based verifier is nearly identical to that of our baseline, SimpleRL-Zoo (Zeng et al., 2025a), which uses training data that is 10 times smaller and less challenging. However, after integrating the model-based verifier, there is a substantial improvement in performance.
Training Efficiency.
Compared to a rule-based verifier, the hybrid verifier incorporates a generative language model into the verification process, increasing verification time during training. However, the HybridEngine’s design (detailed in Appendix E) allows for a substantial increase in GPU utilization, keeping the overhead at an acceptable level. Moreover, the design of our hybrid verifier allows the model-based verifier’s workload to be further reduced.
Evaluation Results using GPT-4o.
Since we have identified the issue of inaccurate judgments by rule-based verifiers, relying on them at inference time – as is common in most previous works – may disadvantage RL training when model-based verifiers are used. Therefore, we also report evaluation results using GPT-4o as the verifier.. As shown in Table 3, most results are similar to those obtained using the rule-based verifier (Table 2), and the conclusions drawn from the latter remain valid. However, a notable exception occurs with the Minerva Math benchmark, where a significant performance gap of 13 points is observed on the Qwen2.5-7B base model when GPT-4o is used as the evaluation verifier (Table 2 vs. Table 3). This suggests that the limitations of rule-based verifiers extend even to well-established benchmarks, highlighting the need for more sophisticated verifiers in certain cases to ensure accurate evaluation.
Table 3: Detailed performance of models across multiple benchmarks with GPT-4o as the verifier. This table evaluates the correctness of the models’ responses by using GPT-4o as the verifier, reporting the best result from each run. Blue lines indicate models trained with a hybrid verifier without evidence of reward hacking, while pink lines indicate runs where reward hacking is detected. “HF” represents HuggingFace Math Verifier.
Model GSM8K
MATH
500
Minerva
Math
Olympiad
Bench
AIME24
(Avg@32)
AMC23 Avg.
Qwen2.5-7B 88.5 65.8 39.0 31.7 0.3 30.0 42.6
      $\hookrightarrow$ + DeepscaleR & HF verifier 93.1 80.8 53.3 45.6 15.3 62.5 58.4
          $\hookrightarrow$ + DS-R1-Distill-Qwen-1.5B verifier 93.7 82.8 50.0 45.6 20.4 70.0 60.4
          $\hookrightarrow$ + R1-Distill-Verifier-1.5B verifier 93.3 80.6 47.1 43.1 17.9 77.5 59.9
          $\hookrightarrow$ + general-verifier 92.7 82.8 52.2 44.1 18.4 70.0 60.0

### v2

# From Accuracy to Robustness: A Study of Rule- and Model-based Verifiers in Mathematical Reasoning

@@ Abstract
Trustworthy verifiers are essential for the success of reinforcement learning with verifiable reward (RLVR), which is the core methodology behind various large reasoning models such as DeepSeek-R1. In complex domains like mathematical reasoning, rule-based verifiers have been widely adopted in previous works to train strong reasoning models. However, the reliability of these verifiers and their impact on the RL training process remain poorly understood. In this work, we take mathematical reasoning as a case study and conduct a comprehensive analysis of various verifiers in both static evaluation and RL training scenarios. First, we find that current open-source rule-based verifiers often fail to recognize equivalent answers presented in different formats across multiple commonly used mathematical datasets, resulting in non-negligible false negative rates. This limitation adversely affects RL training performance and becomes more pronounced as the policy model gets stronger. Subsequently, we investigate model-based verifiers as a potential solution to address these limitations. While the static evaluation shows that model-based verifiers achieve significantly higher verification accuracy, further analysis and RL results imply that they are highly susceptible to hacking, where they misclassify certain patterns in responses as correct, particularly after fine-tuning. This vulnerability is exploited during policy model optimization, leading to artificially inflated rewards. Our findings underscore the unique challenges inherent to both rule-based and model-based verifiers and provide insights toward developing more accurate and robust reward systems for reinforcement learning.
@@ 1 Introduction
Figure 1: The training and evaluation curves of RL on Qwen-2.5-7B using different verifiers, with the x-axis representing training iterations in all plots. Left illustrates the evaluation accuracy averaged over multiple benchmarks, including GSM8K, MATH500, Minerva Math, OlympiadBench, AIME24, and AMC23. Right depicts changes in reward values during training. The “training rewards” indicate the rewards provided by the corresponding reward system to the policy model, whereas the “oracle rewards” represent rewards the model receives when judged by combining with GPT-4o. All benchmarks are reported with a single sample due to computational constraints; detailed stable results at the peak point are provided in Table 2.
In our subsequent RL training experiments, however, we observe that model-based verifiers introduce unique challenges and yield mixed outcomes: while some verifiers can improve RL results by an average of 2.3 absolute points over rule-based verifiers, others are vulnerable to hacking, leading to suboptimal results of RL training(see Figure 1 and Table 2). Reward hacking – a well-known issue in RL – refers to the exploitation of specific patterns by the policy model to deceive the reward system and obtain artificially high rewards (illustrated in the bottom right of Figure 1). Notably, we find that although some model-based verifiers trained on labeled classification data achieve higher classification accuracy than off-the-shelf alternatives, they are more susceptible to hacking during RL training. And we further observe similar phenomena in the general science domain. These findings indicate that the classification accuracy of a verifier does not necessarily reflect its resistance to reward hacking, and therefore may not be a reliable indicator of its effectiveness in RL training.
In the final part of our study, we conduct a systematic probing study into specific hacking patterns that can exploit vulnerabilities in verifiers. We construct a range of adversarial patterns inspired by our case studies, such as the insertion of empty characters or garbled text. Using these constructed “hacking data”, we evaluate whether various model-based verifiers can be deceived. Our results show that all generative verifiers, no matter whether they are specifically fine-tuned for verification, are easily fooled by these patterns. Interestingly, the discriminative verifiers are more robust than the generative ones without the reasoning process.
Our findings in this work clearly underscore the challenges inherent to both rule-based and model-based verifiers primarily in the context of mathematical reasoning: current rule-based verifiers are not sufficiently accurate even for widely used open-source mathematical datasets with short answers that should be easily verifiable. Pursuing more accurate model-based verifiers by fine-tuning is a promising direction to improve RL performance; however, this approach potentially introduces unique vulnerabilities to hacking, which require further investigation in future work.
@@ 3.1 Evaluation Dataset Construction
We curate dataset as a static classification task to examine the capabilities of verifiers in classifying the correctness of model responses with respect to a provided ground-truth answer. The curation process involves three main steps: First, we select and sample from four mathematical RL datasets – Math (Hendrycks et al., 2021), DeepscaleR (Luo et al., 2025), Open-Reasoner-Zero (ORZ-Math)(Hu et al., 2025), and Skywork-OR1(He et al., 2025) – with 1,000 queries sampled from each dataset. In the second step, we generate two responses for each of these queries using two types of language models: (1) Short-CoT models, specifically Qwen2.5-Math-7B-Instruct (Yang et al., 2024b) and Qwen2.5-32B-Instruct (Yang et al., 2024a), and (2) R1-style long CoT models, namely Deepseek-R1-Distill-Qwen-7B and 32B (DeepSeek-AI et al., 2025). Finally, we employ GPT-4o (Hurst et al., 2024) as an annotator to provide ground-truth annotations based on the response and target answer, on whether the model’s response aligns with the target answer, based on a prompt shown in Figure 4 in Appendix B. We further validate GPT-4o’s annotations against human judgments (Appendix B). The final dataset comprises 2,000 examples per dataset, for a total of 8,000 examples. We emphasize that the datasets we selected already represent a relatively easy setting for verification – these datasets contain only short answers, and most were specifically curated to be easily verifiable by rules in order to facilitate RL. Consequently, more realistic scenarios are likely to present greater challenges than those reflected in our empirical results next.
@@ 3.2 Rule-based Verifiers: Precision at the Cost of Recall
We adopt three popular rule-based verifier implementations including: (1) Verl Math Verifier, [^1] https://github.com/volcengine/verl (2) Qwen-Math Verifier, [^2] https://github.com/QwenLM/Qwen2.5-Math and (3) HuggingFace Math Verifier, [^3] https://github.com/huggingface/Math-Verify, following prior work (Zeng et al., 2025b; a; He et al., 2025; Yu et al., 2025). Further implementation details are in Appendix C.
@@
To evaluate the performance of these verifiers, we test them on a subset of data sampled from Deepseek-R1-Distill-Qwen-32B, a state-of-the-art open-source model known for its exceptional mathematical reasoning abilities. As shown in Table 4 in Appendix D, all three verifiers exhibit near-perfect precision ($99\%$). This means that if an answer passes the rules, it is almost certainly correct because the rule-based verifiers rely on deterministic programming language logic and computation. Notably, the HuggingFace Math Verifier and Qwen-Math Verifier show very similar performance. However, the rigid structure of these rule-based systems leads to poor recall, dropping to 0.78 on challenging datasets like Skywork-OR1 (Figure 3). This indicates that there are some correct responses that are misjudged as incorrect, and we illustrate some cases in Figure 5.
@@
As shown in Figure 3, as the capabilities of the models increase, providing accurate supervision becomes more challenging for rule-based verifiers. For example, the recall rate for the Long-CoT models, such as DeepSeek-R1-Distill-Qwen-7B and 32B averages around 0.92, which is much lower than other weaker models. This is because some complex queries, which only advanced models can solve, are misjudged by the rule-based verifier. The inability of rule-based verifiers underlines the difficulty in verifying highly capable models. This trend is particularly concerning, given that the community is advancing increasingly powerful reasoning models, which in turn require stronger verifiers.
@@
Figure 3 shows that as datasets grow more complex, recall rates decline. The Math dataset, simple and well-structured, yields relatively high recall, while harder datasets like Skywork-OR1 show much lower rates. Beyond mathematics, in Appendix J we further analyze the WebInstruct-Verified (Ma et al., 2025) dataset, which spans a broader general science domain. Here, rule-based verifiers perform even worse, with recall dropping below 0.6, highlighting their limited adaptability to diverse and less structured answer formats. These findings underscore a critical limitation: As datasets become more varied, and more challenging, the reliability of rule-based verifiers as supervision tools for scalable reinforcement learning diminishes.
Table 1: Performance of model-based verifiers across datasets, reported as Precision/Recall. To assess them within a hybrid verifier framework, we evaluate samples from DeepSeek-R1-Distill-Qwen-32B, excluding cases already verified correct by HuggingFace Math Verifier (hence N/A). “DS” denotes DeepSeek, and for Qwen series models, the “instruct” suffix is omitted for clarity.
@@ 3.3 Model-based Verifiers: Toward Greater Flexibility
To mitigate the limitation of rule-based verifiers, we next investigate model-based verifiers as a potential alternative. Model-based verifiers seek to leverage the core capabilities of LLMs, including their advanced reasoning skills, to produce more accurate judgments. They are, in principle, better equipped to evaluate answers presented in diverse formats. Model-based verifiers are explored in several concurrent works (Su et al., 2025; Ma et al., 2025; Seed et al., 2025) without deep discussion or ablation on their strengths and limitations. In this section, we first explore model-based verifiers in static evaluation, and in §4 we will discuss its effect in RL training.
@@
We evaluate two categories of general LLM as a verifier: (1) Short-CoT models: Qwen2.5-instruct (1.5B and 7B) (Yang et al., 2024a) and Qwen2.5-Math-instruct (1.5B and 7B) (Yang et al., 2024b). (2) R1-style long-CoT models: DeepSeek-R1-Distill-Qwen (1.5B and 7B) (DeepSeek-AI et al., 2025). We will also discuss model-based verifiers specifically trained for verification tasks in §5. Note that we focus on models with up to 7B parameters, as larger models are neither practical nor efficient for scaling RL training. We note that all these models are generative which will typically generate reasoning traces along with the final judgment. Since rule-based verifiers achieve nearly perfect precision but tend to produce false negatives, we focus here exclusively on the examples that rule-based verifiers classify as incorrect. This approach is able to better distinguish different model-based verifiers. It also aligns with the design of our hybrid verification system in the RL experiments, where rule-based verifiers are applied first, and model-based verifiers are used only for those cases deemed incorrect. We will provide further details in §4.1. Specifically, for the evaluation dataset, we use the subset sampled from DeepSeek-R1-Distill-Qwen-32B, excluding examples that have already been classified as correct by the HuggingFace Math Verifier. For additional details about the evaluation procedure, please refer to Appendix E.
@@ 4 The Effect of Verifiers on RL Training
In §3, we showed that model-based verifiers achieve strong performance across datasets and substantially improve recall on the verification task. Building on this, we adopt model-based verifiers in RL training and compare their impact with rule-based verifiers. Specifically, we utilize the hybrid verifier that integrates the strengths of both approaches. We first evaluate its performance in static settings, then analyze its improvements over rule-based verifiers in RL training.
@@ 4.1 The Hybrid Verifier
In Table 5 in Appendix F, we present the static evaluation results of rule-based, model-based, and hybrid verifiers. The hybrid verifier improves recall by $\sim$3 points over rule-based while maintaining $98\%$ precision. Model-based verifiers alone may exhibit lower recall than the hybrid approach, as smaller models can overthink some straightforward cases. However, integrating the rule-based verifier mitigates this issue, resulting in superior overall performance. In general, the hybrid system achieves superior performance in both precision and recall. Furthermore, by filtering out straightforward cases to the rule-based verifier, the hybrid design substantially reduces the computational load on the model-based verifier. We discuss this further in Appendix G.
@@ 4.2 Experimental Setup
For all experiments, we follow the approach of Deepseek-R1 (DeepSeek-AI et al., 2025), using GRPO (Shao et al., 2024) as the training algorithm and adhering to the zero RL training recipe – starting training directly from the base model. Our policy model is Qwen2.5-7B Base (Yang et al., 2024a), chosen for its practical balance between performance and computational cost, and its widespread use in prior studies (Zeng et al., 2025a; Liu et al., 2025b). We primarily conduct training on the DeepscaleR dataset, owing to its early adoption, high quality, and extensive use in recent work (Liu et al., 2025a; Qu et al., 2025; Aggarwal & Welleck, 2025). To construct a hybrid verifier, we combine the HuggingFace Math Verifier with DeepSeek-R1-Distill-Qwen-1.5B, which achieves the strongest performance among 1.5B-scale models on DeepscaleR (see Table 1). Additional training and evaluation details are provided in Appendix G.
@@
Our evaluation script is based on Yang et al. (2024b), which uses a rule-based verifier. We evaluate on standard mathematical reasoning benchmarks, including GSM8K (Cobbe et al., 2021), MATH 500 (Hendrycks et al., 2021), OlympiadBench (He et al., 2024), and Minerva Math (Lewkowycz et al., 2022), as well as on competition-level benchmarks such as AIME 2024 and AMC 2023. For AIME 2024 and AMC 2023, we report stable results by averaging over 32 random samplings (Avg@32) in Table 2. Further details are provided in Appendix G.
@@ 4.3 Results
Table 2: Detailed performance of models across benchmarks. The best result from each run is reported. Blue lines indicate models trained with a hybrid verifier without evidence of reward hacking, while pink lines indicate runs where reward hacking is detected. “HF” represents HuggingFace Math Verifier. Training and evaluation curves for these models are presented in Figure 1 and Figure 8.
@@
AMC23
(Avg@32)
Avg.
Qwen2.5-7B-SimpleRL-Zoo 91.7 78.2 38.6 40.4 15.6 54.9 53.2
Qwen2.5-7B 88.2 64.6 25.7 30.1 0.3 36.9 41.0
      $\hookrightarrow$ + DeepscaleR & HF verifier 92.8 80.0 37.5 42.2 15.3 62.3 55.0
          $\hookrightarrow$ + DS-R1-Distill-Qwen-1.5B verifier 93.3 82.4 41.2 42.5 20.4 64.1 57.3
          $\hookrightarrow$ + R1-Distill-Verifier-1.5B verifier 93.0 79.8 40.4 40.1 17.8 62.2 55.6
          $\hookrightarrow$ + general-verifier 92.5 82.0 43.0 40.9 18.4 65.2 57.0
@@
As shown in Figure 1 and Table 2, incorporating the hybrid verifier yields a substantial improvement in evaluation accuracy, reaching a peak of 57.3 – 2.3 points higher than using the rule-based verifier alone. In addition, the hybrid verifier enhances dataset utilization by reducing the fraction of responses that cannot be successfully parsed. For example, Table 2 shows that the performance of the rule-based verifier is only marginally better than our baseline, SimpleRL-Zoo (Zeng et al., 2025a), which uses training data that is 10 times smaller and less challenging. By contrast, integrating a model-based verifier leads to a more pronounced improvement in overall performance.
Cross-Dataset Generalization.
To further test the generalization of our findings, we conduct RL experiments on the Skywork-OR1 dataset (He et al., 2025), as reported in Table 6 in Appendix I. The results confirm that the limitations of rule-based verifiers also persist in this setting, with a clear performance gap between using only a rule-based verifier and incorporating the hybrid verifier. The improvements are substantial across multiple benchmarks, including +3.6 points on MATH500, +4.0 points on OlympiadBench, and +6.6 points on AMC23.